Clculate the accelerating potential of a proton beam to be given a effective wavelength of 0.005A°

Grade:11

1 Answers

Vikas TU
6 years ago
Frm the de-brogliew wavelength already Potential and wavelength formuae derived for the charge like electron or proton,
we know that,
wavelength k = root(150/V)
Or
k = 12.3/root(V)
Putting the values we get,
root(V) = 12.3/(0.005*10^-10) = > 2.46*10^13
Squaring both sides we get,
V = 6.05*10^26 Volts.